Drive the following formulas;

[xvYv(x)]=xvYv-1(x)


1
Expert's answer
2021-03-09T05:59:52-0500

Taking LHS

"\dfrac{d}{dx}[x^vY_v(x)]"


Now let "m=Y_v(x),n=x^v"


Now using Lebinitz theorem we have-

"\dfrac{d}{dx}[x^vY_v(x)]"


"=^1C_0(1! \times x^v \times Y_{v-1}(x))"


"=x^vY_{v-1}(x)"